OpenStudy (anonymous):

A Go Cart is bought for $4200 and sold 5 years later for $1130. Assuming the function is linear, find the depreciation formula and the value of the Go Cart after 3 years.

OpenStudy (anonymous):

were assuming it's linear, therefore the equation should look like: y = mx + b assuming it's linear also implies that the slope is constant for all values. m = (y2 - y1)/(x2 - x1) = (1130 - 4200)/(5 - 0) m = -614 in y = mx + b, b is the initial value of y, and y is really a function of x [f(x)], varying with x. y = -614x + 4200 is the equation you now use, where y is the value of the go-cart for x years. (notice x = 0, gives 4200, which makes sense). plugging x = 3 gives you your answer for 3 years.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

glad i could help :)